OpenStudy (anonymous):

Find the area of the circle in simplest polynomial form. (A = 3.14r2) If the radius of the circle is 2x + 1.

OpenStudy (anonymous):

\[A=\pi(2x+1)^2\]

OpenStudy (anonymous):

Step by step, like a gee. (2x+1)(2x+1)=4x^2+4x+1 <--simplest polynomial form (3.14)(4x^2+4x+1)=1384.74 Area=1384.74 ^_^

OpenStudy (anonymous):

\[\pi(4x^2+4x+1)\] \[4\pi x^2+4\pi x+\pi\]